Solve the following system of linear equations by the elimination method. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)
5/12x-1/2y=-61/6
1/5x-1/3y=-167/15
x=?
y=?
Please show me the steps to get the answer. Thanks

Answer by mananth(16946)

5/12x-1/2y=-61/6
LCD=12. Multiply equation bty 12
5x-6y=-122.....................1
1/5x-1/3y=-167/15
LCD = 15
3x-5y=-167......................2
5 x -6 y = -122 .............1
3 x -5 y = -167 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 6
-25 x + 30 y = 610
18 x + -30 y = -1002
Add the two equations
-7x=-392
/ -7
x = 56 = 56.00
plug value of x in (1)
5 x + -6 y = -122
280 + -6 y = -122
-6 y = -122-280
-6 y = -402
y = 67